The points (-2,11) and (6,3) lie on the same line. What is the x-intercept of the line?
Answer by rfer(16322)

m=3-11/6--2
m=-7/8
----------------
y=mx+b
3=-7/8*6+b
3=-42/6+b
3=-7+b
b=10
-----------------
y=-7/8x+10
0-10=-7/8x
x=11.43
(11.43,0)